Circular No. 8361

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2004cs, 2004ct, AND 2004cu Three apparent supernovae have been discovered on unfiltered CCD

images: SNe 2004cs and 2004cu by LOSS/KAIT (via W. Li and D. Singer; cf. IAUC 8357), and SN 2004ct by T. Boles (cf. IAUC 8343).

SN  2004 UT  α_{2000}  δ_{2000}  Mag.  Offset
2004cs June 23.42 17^{h}50^{m}14.^{s}38 +14^{o}16′59′′.5 18.0 10′′.4 E, 15′′.2 S
2004ct June 22.01 15 43 45.75 +28 24 59.5 17.5 4′′.1 W, 4′′.4 N
2004cu June 24.28 14 18 29.82 +12 53 03.0 18.8 27′′.1 E, 3′′.5 N

Additional approximate magnitudes, from images by the respective discoverers unless otherwise noted: SN 2004cs in UGC 11001, June 21.36 UT, [19.0; 24.20, 17.8. SN 2004ct in MCG +05-37-17, Mar. 20, [19.5; May 16, [19.5; 24.201, 17.5 (KAIT). SN 2004cu in NGC 5550, May 17.36, [20.0; June 12.33, 19.3 (marginal detection). Boles adds that SN 2004ct is not present on Digitized Sky Survey red (1989) or blue (1993) plates.

COMET C/2003 K4 (LINEAR) Further to IAUC 8358, R. W. Russell, D. L. Kim, M. L. Sitko, and

W. J. Carpenter report 3–13-μm spectrophotometry of comet C/2003 K4, obtained on June 20.3 UT at Mt. Lemmon (integration times 90 min on the comet; reference star α Boo): “A continuum, smooth to within the signal-to-noise, was seen to rise from 3.5 to 8 μm, beyond which a weak silicate emission band may have been observed. An underlying blackbody continuum with a temperature of ∼ 235 ± 10 K was fit to the continuum fluxes at 8.4 and 12 μm. This grain temperature is ∼ 22 ± 5 percent higher than that of an equilibrium blackbody at the comet’s heliocentric distance. Using the same wavelength region (10.34–10.71 μm) as for other, brighter comets in order to calculate a silicate-feature-to-continuum ratio, the possible silicate feature was ∼ 1.10 ± 0.05 times higher than the continuum, with the silicate-feature-to-continuum ratio > 1. The comet showed the following narrowband (∼ 0.25 μm) magnitudes and combined random errors: [8.0 μm] = 4.25 ± 0.10; [10.5 μm] = 2.41 ± 0.06; [12 μm] = 1.76 ± 0.06. Due to the low flux level of the comet and the weakness of its silicate feature, no structure due to crystalline material was discernible.”

Visual total-magnitude estimates: May 12.96 UT, 10.6 (A. Baransky, Bucha, Ukraine, 0.20-m refl.); 24.98, 9.8 (A. Diepvens, Balen, Belgium, 0.15-m refractor); June 7.96, 8.8 (K. Hornoch, Lelekovice, Czech Rep., 10×80 binoc.); 21.93, 8.0 (J. J. Gonzalez, Asturias, Spain, 7×50 binoc.).

2004 June 24  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green